OpenStudy (anonymous):

A fishing boat plots its location at the coordinates (110, 0) on a map. A school of tuna is located at (125, -20). What is the straight line distance from the boat to the tuna? a. 30 b. 15 c. 10 d. 25

OpenStudy (anonymous):

D

OpenStudy (anonymous):

110-125=-15 0-20=20 sqrt((-15)^2+20^2)=25